OpenStudy (anonymous):

Use the Rational Zeros Theorem to write a list of all possible rational zeros of the function. f(x) = -2x4 + 4x3 + 3x2 + 18

OpenStudy (anonymous):

Please Help ASAP!!

OpenStudy (amistre64):

use the thrm ... to list the possible ... it tells you exactly what to do :/

OpenStudy (anonymous):

But I don't know how to do it

OpenStudy (amistre64):

tell me what the thrm states ...

OpenStudy (anonymous):

If P(x) is a polynomial with integer coefficients and if is a zero of P(x) ( P() = 0 ), then p is a factor of the constant term of P(x) and q is a factor of the leading coefficient of P(x) .

OpenStudy (amistre64):

good ... seems a little incomplete but lets say it like this: we have a poly of the form:\[P(x)=ax^n+bx^{n-1}+...+jx+k\] IF there exists a rational root to this poly, then it will have to be of the form:\[\pm\frac{factors~of~k}{factors~of~a}\]

OpenStudy (amistre64):

what are the factors of the constant: 18 what are the factors of leading coeff: 2

OpenStudy (anonymous):

Factors of 2 are: 2*1 Factors of 18 are:18*1, 2*9, 3*6,

OpenStudy (amistre64):

very good now lets see how we can pair them up:\[\pm\frac{1,2,3,6,9,18}{1,2}\] \[\pm\frac{1}{1},\pm\frac{2}{1},\pm\frac{3}{1},...,\pm\frac{18}{1}\] \[\pm\frac{1}{2},\pm\frac{2}{2},\pm\frac{3}{2},...,\pm\frac{18}{2}\] reduce and remove duplicates to simplify the list.

OpenStudy (amistre64):

show me your final list, or ask me questions if its unclear :)

OpenStudy (anonymous):

you can just remove 2/1 and 2/2 since they are the same

OpenStudy (amistre64):

2/1 does not equal 2/2 .... i believe what your thought was: 1/1 = 2/2: 1=1. having two 1s in the list is redundant.

OpenStudy (amistre64):

2/2, 6/2, and 18/2 create duplicate elements, but the other /2 elements are part of the list if i see it right

OpenStudy (anonymous):

Ok

OpenStudy (amistre64):

to compare, that leaves us: +- (1,2,3,6,9,18, 1/2, 3/2, 9/2) as all the potential rational roots

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

does it make more sense now?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

good luck :)